It took a cyclist 6 hours to travel 48 miles going against the wind. The next day on the return trip, it took the cyclist 3 hours traveling with the wind. What was the speed of the cyclist?
Answer by mananth(16946)

speed of cyclist = x
speed of wind = y
with wind speed = x+y
against wind speed = x-y
d=rt
d is same in this case
r=d/t
x-y = 180/6
x-y=30
x+y=180/3
x+y=60
solve for x & y
2x=90
x=45
y=15